# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 6.4 Security Analysis of Profile P2
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model against static malicious adversaries corrupting up to N1N-1 parties, with security-with-abort. Concretely:
$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$ with $m=nk=1536$ coefficients and $\epsilon_{\mathsf{edaBits}}<2^{-12}$ with cut-and-choose parameter $B=64$.
Proof 6.3.
@@
    $\mathsf{H}_{1}$: Replace SPDZ with $\mathcal{F}_{\mathsf{SPDZ}}$ (loss: $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$)
@@
    $\mathsf{H}_{5}$: Backward-simulate AND tree given result bit (loss: 0)
@@ 6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
Profile P3+ designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC. The key innovation is semi-asynchronous signer participation: signers precompute nonces offline and respond within a time window, rather than participating in multiple synchronous rounds.
@@ 6.5.1 UC Security Analysis of Profile P3+
  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
@@
$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}}^{\mathsf{2PC}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$
where $\kappa$ is the computational security parameter.
Proof 6.6.
@@ Conclusion.
Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$.
@@
    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.4); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
@@
P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes $\approx$30% per-attempt success rate (3.7 attempts for P3+, 3.0 for P2).
@@ 6.7 Additional Extensions
Pairwise seeds can be established via ML-KEM [FIPS203] for full post-quantum security with forward secrecy.
@@ 7.1 Positioning Among Concurrent Work
Concurrent work includes Bienstock et al. [BDEPT25] (arbitrary $T$, UC security, but honest-majority and 37–136 rounds), Celi et al. [CDENP26] (dishonest-majority, 6 rounds, but $T\leq 6$), and Trilithium [Trilithium25] (UC-secure, but 2-party only). Our Profile P2 targets a different point in the design space: arbitrary $T$ with dishonest-majority security in 5 constant rounds.
@@ 7.2 Summary of Contributions
Our primary technical contribution is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing that structurally matches the long-term secret. Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.4). The bounded-nonce constraint prevents perfect uniformity (Remark 4.6), but the min-entropy guarantee is more than sufficient for practical security. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy: signing sets of size $|S|\geq T$ suffice, matching the information-theoretic minimum for Shamir reconstruction. Individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator (not broadcast), preventing a key-recovery attack via the injective matrix $\mathbf{A}$ (Remark 4.1). Commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model.
As a secondary technique, we introduce pairwise-canceling masks adapted from ECDSA [CGGMP20] to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$. This addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security.
@@
Our optimized Rust implementation demonstrates practical efficiency: P1 and P3+ achieve 4–94 ms for configurations from $(2,3)$ to $(32,45)$ threshold; P2 ranges from 21–194 ms, trading latency for eliminating hardware trust. Success rates of $\approx$21–45% are comparable to single-signer ML-DSA. This represents a $4.8\times 10^{10}$ improvement over naive threshold approaches for $T=16$. All signatures are syntactically identical to single-signer ML-DSA and pass standard FIPS 204 verification without modification.
@@ 7.3 Limitations and Future Work
Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds. The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.8 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing. The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the per-coordinate Rényi divergence is small ($R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$), the full-vector composition $(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ grows to $\approx 2^{5.4}$ for $|S|=17$ ($\approx$3-bit loss) and $\approx 2^{75}$ for $|S|=33$ ($\approx$38-bit loss); see Theorem 4.11. For the pairwise masks used on commitment and r0-check values, $|S\setminus C|\geq 2$ honest parties are needed. In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for response privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values. Finally, we assume a synchronous model where all parties are online simultaneously during signing (though P3+ relaxes this to semi-asynchronous participation).
@@
The most natural theoretical extension is adaptive security. Our analysis assumes a static adversary that selects which parties to corrupt before protocol execution begins; extending the results to the adaptive setting—where the adversary may corrupt parties during the protocol based on observed transcripts—remains open. The Shamir nonce DKG introduces a subtle obstacle: each honest party’s polynomial $\mathbf{f}_{h}$ has its leading coefficient $\mathbf{a}_{h,T-1}$ sampled uniformly, but the remaining coefficients are constrained by the constant term $\hat{\mathbf{y}}_{h}$ and the evaluation points. Under adaptive corruption, the simulator must equivocate these polynomials after the fact, which may require an erasure model (parties delete intermediate state after each round) or a non-committing encryption wrapper around the DKG messages. The analogous question for ECDSA threshold signatures was resolved by Canetti et al. [CGGMP20] using UC-secure commitments; whether a similar approach suffices in the lattice setting, where the polynomial structure imposes additional algebraic constraints, is unclear.
@@
On the protocol side, three open problems stand out. The most practically significant concerns eliminating the Irwin-Hall security gap for large signing sets. Our proven security bound degrades with $|S|$ because the aggregated nonce $\mathbf{y}=\sum_{i}\lambda_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than the uniform distribution assumed by FIPS 204: the full-vector Rényi divergence $(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ grows from $\approx 2^{5.4}$ ($|S|=17$, $\approx$3-bit security loss) to $\approx 2^{75}$ ($|S|=33$, $\approx$38-bit loss). For practical deployments with $|S|\leq 17$—covering the vast majority of HSM clusters, custody quorums, and certificate authority configurations—the proven bound provides full 128-bit security. For larger committees, the distribution-preserving rejection sampling technique of Raccoon [Raccoon2024] offers a principled solution: after computing the aggregated nonce, parties collectively evaluate a rejection function that flattens the Irwin-Hall distribution to uniform, yielding $R_{2}^{\mathsf{vec}}=1$ (no security loss) at the cost of an additional interactive rejection round and a reduced per-attempt success rate. The key technical challenge in the threshold setting is that $\mathbf{y}$ is secret-shared—no single party knows the full nonce—so the rejection decision must be computed via MPC over the shares, adding communication complexity proportional to the rejection function’s circuit size. Whether a lightweight, non-interactive variant exists (e.g., each party independently rejecting based on local share statistics) remains open.
Second, this work focuses exclusively on ML-DSA-65 (NIST Security Level 3). The other two standardized parameter sets present distinct challenges: ML-DSA-44 (Level 2) has a smaller masking range $\gamma_{1}=2^{17}$, which tightens the per-coordinate Rényi bound by a factor of $\approx 16$, potentially pushing $R_{2}^{\epsilon,\mathsf{coord}}$ above $1.05$ for moderate signing sets and making the full-vector composition correspondingly more severe. Conversely, ML-DSA-87 (Level 5) uses $k=8,\ell=7$, increasing the r0-check circuit from 1536 to 2048 coefficients and the nonce DKG message size from 3.8 KB to 7.2 KB. A systematic parameter study with benchmarks across all three levels would be valuable for deployment planning.
@@
Third, in P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.9). In the extreme case $T=N$ (unanimous signing), a single honest party’s commitment and r0-check values are not hidden in P2. (In P1 and P3+, the coordinator-only communication model protects $\mathbf{W}_{i}$ even with $|S\setminus C|=1$; see Remark 4.1.) Replacing pairwise PRF masks with a non-interactive zero-sharing scheme—where each party’s mask is derived from a common reference string rather than pairwise seeds—or employing a verifiable PRF construction could close this gap, yielding uniform $|S|\geq T$ privacy across all revealed values. Additionally, achieving true information-theoretic privacy (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.6), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$.

### v5

@@ 6.4 Security Analysis of Profile P2
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ against static malicious adversaries (corrupting up to N1N-1 parties) in the $(\mathcal{F}_{\mathsf{SPDZ}},\allowbreak\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model, with security-with-abort. Concretely:
$\bigl|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]\bigr|$
$\quad\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq m/q<2^{-12}$ with $m=nk=1536$ coefficients ($1536/8380417\approx 2^{-12.4}$), and $\epsilon_{\mathsf{edaBits}}\leq m/q+2^{-B}<2^{-12}$ where $B=64$ makes the cut-and-choose term $2^{-64}$ negligible (the dominant term is again $m/q$ from the internal SPDZ MAC check within edaBits generation).
Proof 6.4.
@@
    $\mathsf{H}_{1}$: Replace SPDZ with $\mathcal{F}_{\mathsf{SPDZ}}$ (loss: $\epsilon_{\mathsf{SPDZ}}\leq m/q<2^{-12}$)
@@
    $\mathsf{H}_{5}$: Backward-simulate AND tree given result bit: for each comparison bit $b_{j}$, the simulator samples honest additive shares uniformly subject to summing to the known $b_{j}$ value; SPDZ additive shares are uniform conditioned on their sum, so this is statistically identical to the real distribution (loss: 0)
@@ 6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
Profile P3+ designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC. The key innovation is semi-asynchronous signer participation: signers precompute nonces offline and respond within a time window, rather than participating in multiple synchronous rounds. We use the term semi-asynchronous to denote that signers need not be online simultaneously for interactive rounds, but must respond within a bounded validity window $\Delta$; the 2PC subprotocol between the two CPs executes synchronously independently of the signers.
@@ 6.5.1 UC Security Analysis of Profile P3+
  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$, where $\mathbf{w}=\mathbf{A}\mathbf{y}$ is the full (unmasked) commitment vector (not just $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$; $\mathbf{w}$ is passed as $\mathsf{CP}_{1}$’s private input by the combiner), and the private shares satisfy $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}\bmod q$ (arbitrary additive split; in the concrete protocol, $\mathbf{S}_{j}=\sum_{i\text{ assigned to }\mathsf{CP}_{j}}\lambda_{i}\cdot c\mathbf{s}_{2,i}$)
@@
$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}}^{\mathsf{2PC}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$
where $\kappa$ is the computational security parameter and $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the LP07 cut-and-choose extraction error when $\mathsf{CP}_{1}$ is corrupted.
Proof 6.7.
@@ Conclusion.
Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$ (where $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ from LP07 cut-and-choose extraction).
@@
    Privacy: Nonce shares $\mathbf{y}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
Proof 6.9.
(1) UC security follows from Theorem 6.6 via the UC composition theorem [Canetti01]. (2) EUF-CMA security: by Lemma 3.6, masks cancel in the aggregate, so $(\tilde{c},\mathbf{z},\mathbf{h})$ has identical structure to single-signer ML-DSA; Theorem 4.2 applies. (3) Privacy: nonce shares $\mathbf{y}_{i}$ have statistical privacy via Theorem 4.7; key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; Remark 4.9—Theorem 4.7 establishes only nonce share privacy, not key privacy from $\sigma$); masked values $\mathbf{W}_{i}$, $\mathbf{V}_{i}$ are sent to the combiner (trusted under 1-of-2 CP honest assumption) or protected by mask hiding (Lemma 4.15) when $|S\setminus C|\geq 2$.
@@
P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes 27–33% per-attempt success rate depending on configuration (3.7 attempts for P3+, 3.0 for P2).
@@ 6.7 Additional Extensions
Pairwise seeds can be established via ML-KEM [FIPS203] for full post-quantum security (static pairwise seeds; per-session forward secrecy requires ephemeral per-round encapsulation).
@@ 7.1 Positioning Among Concurrent Work
Concurrent work includes Bienstock et al. [BdCE25] (arbitrary $T$, UC security, but honest-majority and 23–79 rounds per [BdCE25]), Celi et al. [CDENP26] (dishonest-majority, 6 rounds, but $T\leq 6$), and Trilithium [Trilithium25] (UC-secure, but 2-party only). Our Profile P2 targets a different point in the design space: arbitrary $T$ with dishonest-majority security in 5 constant rounds.
@@ 7.2 Summary of Contributions
Our primary technical contribution is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing that structurally matches the long-term secret. Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.7). Key privacy given the public signature relies on the same lattice hardness as single-signer ML-DSA (computational, informal; formal reduction is an open problem for ML-DSA-based constructions, unresolved even in the single-signer setting). The bounded-nonce constraint prevents perfect uniformity (Remark 4.10), but the min-entropy guarantee is more than sufficient for practical security. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for nonce share privacy: signing sets of size $|S|\geq T$ suffice, matching the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction. Individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator (not broadcast), preventing a key-recovery attack via the injective matrix $\mathbf{A}$ (Remark 4.1). Commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.15), which applies to P2’s broadcast model.
As a secondary technique, we introduce pairwise-canceling masks (adapted from ECDSA [CGGMP20]) to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$. This addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security. The Irwin-Hall security gap of prior approaches is resolved by direct shift-invariance (Corollary Q.24): the EUF-CMA loss is <0.013<0.013 bits per signing session for $|S|\leq 33$ (total: $0.013\cdot q_{s}$ bits), giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$), remaining below $1$ bit per session for all $|S|\leq 2{,}584$.
@@
Our optimized Rust implementation demonstrates practical efficiency: P1 achieves 4–60 ms and P3+ achieves 17–94 ms for configurations from $(2,3)$ to $(32,45)$ threshold; P2 ranges from 21–194 ms, trading latency for eliminating hardware trust. Success rates of $\approx$21–45% (Table 5) are comparable to single-signer ML-DSA. All signatures are syntactically identical to single-signer ML-DSA and pass standard FIPS 204 verification without modification.
@@ 7.3 Limitations and Future Work
Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds. The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.6 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing. The aggregated nonce follows an Irwin-Hall rather than uniform distribution. By the direct shift-invariance analysis (Corollary Q.24), the per-coordinate chi-squared divergence between real and simulated nonces is $\chi^{2}(\mathsf{IH}+\beta\|\mathsf{IH})<7\times 10^{-6}$ for $|S|\leq 17$ and $1.4\times 10^{-5}$ for $|S|\leq 33$, giving a full-vector EUF-CMA security loss <0.013<0.013 bits per signing session for $|S|\leq 33$ (total: $0.013\cdot q_{s}$ bits)—at most $13$ bits for $q_{s}\leq 1{,}000$ (or <7<7 bits in the $|S|\leq 17$ regime using the <0.007<0.007 per-session bound), well below the 192-bit security target. For the pairwise masks used on commitment and r0-check values, $|S\setminus C|\geq 2$ honest parties are needed. In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for nonce share privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values. This work focuses on ML-DSA-65 (NIST Level 3); extending to ML-DSA-44 ($\gamma_{1}=2^{17}$, $\beta=\tau\eta=78$) and ML-DSA-87 ($k=8,\ell=7$, $\gamma_{1}=2^{19}$, $\beta=120$) is straightforward in principle—ML-DSA-44 loosens the chi-squared bound by ${\approx}2.5\times$ relative to ML-DSA-65 (smaller $\gamma_{1}$ increases the ratio $\beta/\gamma_{1}$) but retains <0.04<0.04 bits of EUF-CMA loss per session for $|S|\leq 33$, while ML-DSA-87’s smaller $\beta$ at the same $\gamma_{1}$ tightens the bound by ${\approx}2.7\times$ (<0.007<0.007 bits per session for $|S|\leq 33$); the latter also expands the r0-check circuit from 1536 to 2048 coefficients; a full benchmark across all three levels remains to be conducted. Finally, we assume a synchronous model where all parties are online simultaneously during signing (though P3+ relaxes this to semi-asynchronous participation).
@@
The most natural theoretical extension is adaptive security. Our analysis assumes a static adversary that selects which parties to corrupt before protocol execution begins; extending the results to the adaptive setting—where the adversary may corrupt parties during the protocol based on observed transcripts—remains open. The Shamir nonce DKG introduces a subtle obstacle: each honest party’s polynomial $\mathbf{f}_{h}$ has its leading coefficient $\mathbf{a}_{h,T-1}$ sampled uniformly, but the remaining coefficients are constrained by the constant term $\hat{\mathbf{y}}_{h}$ and the evaluation points. Under adaptive corruption, the simulator must equivocate these polynomials after the fact, which may require an erasure model (parties delete intermediate state after each round) or a non-committing encryption wrapper around the DKG messages. The analogous question for ECDSA threshold signatures was resolved by Canetti et al. [CGGMP20] using UC-secure commitments; whether a similar approach suffices in the lattice setting, where the polynomial structure imposes additional algebraic constraints, is unclear. Separately, a proof in the quantum random oracle model (QROM) would strengthen the EUF-CMA bound against quantum adversaries making superposition queries to the challenge hash—an upgrade shared with single-signer FIPS 204 [FIPS204].
@@
A formal security reduction for key privacy—the property that a signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ does not reveal $(\mathbf{s}_{1},\mathbf{s}_{2})$ beyond the public key $\mathbf{t}_{1}$—remains open even for single-signer FIPS 204 ML-DSA and is inherited by all threshold variants. In the threshold setting, additional leakage vectors arise: partial responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ and masked commitments $\mathbf{W}_{i}$ are visible to the coordinator or broadcast to the signing group. Theorem 4.7 establishes that nonce shares $\mathbf{y}_{h}$ retain high conditional min-entropy given the adversary’s view, bounding the information about $\mathbf{s}_{1,h}$ leaked by any single partial response; however, a formal reduction showing that the aggregate response $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$ and hint vector $\mathbf{h}$, over polynomially many queries with Irwin-Hall nonces and Shamir share structure, do not expose $\mathbf{s}_{1}$ to a lattice adversary remains to be constructed.
@@
Decentralized Key Generation.
Our construction assumes a trusted dealer for key setup: the dealer samples the short-coefficient signing key $(\mathbf{s}_{1},\mathbf{s}_{2})\leftarrow\chi^{\ell}\times\chi^{k}$, computes the public key $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$, and distributes Shamir shares $\mathbf{s}_{1,i}$, $\mathbf{s}_{2,i}$ directly. Fully decentralized key generation for ML-DSA requires verifiably short-coefficient Shamir sharing: each party must prove $\|\mathbf{s}_{1,h}\|_{\infty}\leq\eta$ without reconstruction. Standard Feldman VSS does not apply since Shamir shares of $\eta$-small secrets lie in $\mathbb{Z}_{q}$ and reveal no information about the coefficient bound; lattice-based zero-knowledge techniques—such as BDLOP commitments with norm proofs [BDLOP18]—offer a path forward. We leave a complete treatment to future work.
On the protocol side, three open problems stand out.
Eliminating offline nonce preprocessing.
The Shamir nonce DKG requires one offline preprocessing round of $O(N^{2})$ messages per signing session, structurally necessary because nonce shares must be committed before the challenge hash is computed to prevent adversarial rewinding. A natural question is whether this phase can be merged into the online protocol: a combined first round where parties simultaneously distribute nonce shares and broadcast the nonce commitment $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\sum_{i}\hat{\mathbf{y}}_{i},2\gamma_{2})$ would eliminate the offline cost at the price of one additional online round. Whether such a per-session DKG preserves statistical nonce privacy against an adversary who aborts after observing first-round messages—and whether the SPDZ preprocessing in Profile P2 can be batched with nonce DKG preprocessing into a single unified offline phase—are open questions with direct impact on deployment latency.
Reducing the MPC frontier via carry separation.
Profile P2’s MPC evaluates the full $r_{0}$-check over $\mathbb{Z}_{q}$ (23-bit arithmetic). A finer decomposition is possible: because each party holds Shamir nonce and key shares, the low-bit residual $\ell_{i}=\mathsf{LowBits}(A\mathbf{y}_{i}-c\mathbf{s}_{2,i})$ is locally precomputable without communication. The $r_{0}$-check then reduces to evaluating only the carry of $\sum\lambda_{i}\ell_{i}$ and a single comparison bit—a much lighter computation than the full modular circuit. With standard Shamir coefficients $\lambda_{i}\approx q$, this yields no circuit reduction, since the intermediate sum overflows $\mathbb{Z}_{q}$. However, if reconstruction coefficients satisfy $\lambda_{i}\in\{0,1\}$—as in Vandermonde-style short-coefficient LSSS explored by concurrent work [BCDENP25]—the carry is bounded by $\lfloor T/2\rfloor$, reducing the MPC computation to a $\lceil\log_{2}T\rceil$-bit carry circuit rather than a full 23-bit modular arithmetic circuit. Designing a security proof for this hybrid (Shamir nonce DKG for statistical nonce privacy combined with short-coefficient sharing for $\mathbf{s}_{2}$) is an open problem; the key challenge is showing that the bounded carry value does not leak sufficient information about $\mathbf{s}_{2}$ to assist a lattice adversary. More broadly, whether threshold ML-DSA with dishonest-majority and arbitrary $T$ can be instantiated without general-purpose MPC—overcoming the construction-specific $T\leq 6$ barrier of existing short-coefficient LSSS schemes (the Ball-Çakan-Malkin bound [BCM21] establishes an asymptotic $\Omega(T\log T)$ share-size lower bound; the specific $T\leq 6$ ceiling is Celi et al.’s construction-specific limit)—remains a fundamental open question in lattice-based threshold cryptography.
Commitment mask hiding for $T=N$.
@@
In P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.15). In the extreme case $T=N$ (unanimous signing), a single honest party’s commitment and r0-check values are not hidden in P2. (In P1 and P3+, the coordinator-only communication model protects $\mathbf{W}_{i}$ even with $|S\setminus C|=1$; see Remark 4.1.) Replacing pairwise PRF masks with a non-interactive zero-sharing scheme—where each party’s mask is derived from a common reference string rather than pairwise seeds—or employing a verifiable PRF construction could close this gap, yielding uniform $|S|\geq T$ privacy across all revealed values. Additionally, achieving perfect uniformity (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.10), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$. (Note: the term “statistical privacy” in this work refers to the high-min-entropy guarantee of Theorem 4.7, not SD =0=0.)